OpenStudy (anonymous):

how many molecules of liquid ethanol are in 10.00 ml of ethanol, given its density is 0.7918 g/mL

OpenStudy (anonymous):

do you have the density formula?

OpenStudy (anonymous):

D = M/V

OpenStudy (anonymous):

so... 0.7918=M/10.00 right?

OpenStudy (anonymous):

Yea that's correct

OpenStudy (anonymous):

then what is the mass in grams?

OpenStudy (anonymous):

.07918 g

OpenStudy (anonymous):

?

OpenStudy (anonymous):

0.7918x10.00=7.918

OpenStudy (anonymous):

Wait why did you multiply?

OpenStudy (anonymous):

reverse pamdas... to get the M alone to solve for it.

OpenStudy (anonymous):

Oh ok i see

OpenStudy (anonymous):

so what is the compound for ethanol?

OpenStudy (anonymous):

I don't understand..sorry

OpenStudy (anonymous):

what is the equation for it? the atoms.

OpenStudy (anonymous):

Do you mean how many grams per mole of ethanol?

OpenStudy (anonymous):

CH(sub 3)OH

OpenStudy (anonymous):

wait in the question i meant methanol sorry

OpenStudy (anonymous):

okay what is the formula mass for it? the sum of masses of the atoms?

OpenStudy (anonymous):

12.10+3(1.01)+16+1.01=?

OpenStudy (anonymous):

thats a 12.01

OpenStudy (anonymous):

12.01 + 1.01*3 + 16 + 1.01

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

32.05

OpenStudy (anonymous):

 conversion chart.

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yea thankyou

OpenStudy (anonymous):

yeah anytime.

OpenStudy (anonymous):

1.48 * 10^23